OpenStudy (anonymous):

The polynomial x^3 + 5x^2 -57x -189 expresses the volume, in cubic inches, of a shipping box, and the width is (x+3) in. If the width of the box is 15 in., what are the other two dimensions? ( Hint: The height is greater than the depth.) A. height: 19 in. depth: 5 in. B. height: 21 in. depth: 5 in. C. height: 19 in. depth: 7 in. D. height: 21 in. depth: 7 in.

OpenStudy (alekos):

need help?

OpenStudy (anonymous):

it is b correct?

OpenStudy (alekos):

that's what I got. do you know how to get this?

OpenStudy (anonymous):

Given: \[v=x^3+5x^2-57x-189\] The volume of a box can be written as \[v=height \times depth \times width\] The width of the box can be written as \[w=x+3=15\] Which means that x is 12 \[x=15-3=12\] Plug that into the polynomial to get the volume \[v=12^3+(5\times12^2)-(57\times12)-189=1728+(5\times144)-684-189\]\[=1728+720-873=1575\] Plug the values into the second volume equation \[1575=h \times d \times 15, h \times d=1575\div15=105\] Multiply the heights and depths of the choices to find which has a product of 105 \[A: 19\times5=95, B: 21\times5=105\]

OpenStudy (alekos):

rebecca. thanks for helping, but it's better to let samantha work it out